# Beyond Pairwise: Global Zero-shot Temporal Graph Generation

### v2

@@ 1 Introduction
The TRE task is formulated as follows: given a pair of event mentions, identify the temporal relation between them. The task has seen significant progress in recent years with the development of supervised models Tan et al. (2023); Niu et al. (2024). However, these models require large amounts of training data, which is scarce in most domains and languages, and difficult to obtain due to the complexity of manually annotating such relations Pustejovsky and Stubbs (2011).
@@
To address the scarcity of fully annotated datasets, and the inefficiency and global inconsistency of pairwise classification in zero-shot settings, we contribute the following. First, we propose a novel zero-shot LLM method that generates the entire temporal graph in a single inference step. Our method prompts the model to produce a free-form summary of the event timeline to guide reasoning or “thinking”, followed by classification of all event pairs, aggregated via a global temporal constraints optimization algorithm to ensure consistency (Figure 1, §4). This approach significantly reduces computational cost compared to pairwise methods while achieving competitive performance (§6). Second, we introduce OmniTemp (will be made publicly available), a new dataset with full temporal relation annotations for all targeted event pairs, supporting global temporal graph extraction (§3). Lastly, we discuss how annotation scope and guideline inconsistencies affect zero-shot model assessment. We show that limiting annotations to short-distance event pairs, as well as discrepancies between widely used datasets such as MATRES and TB-Dense, can hinder fair and reliable evaluation of TRE models in zero-shot settings.
@@ 2.1 Temporal Relation Extraction Datasets
The temporal relation extraction task aims to determine the temporal order between pre-extracted events in a text Pustejovsky et al. (2003). For fair and unbiased model evaluation, datasets should provide gold labels for all event pairs or, at a minimum, be randomly sampled from the full set. However, most existing datasets for temporal relation extraction provide only partial annotation due to the complexity and cost of the process. As a result, the two most widely used datasets, MATRES Ning et al. (2018b) and TimeBank-Dense (TB-Dense) Chambers et al. (2014), annotate only relations between events in consecutive sentences.
@@
While NarrativeTime provides an exhaustively annotated dataset, it follows complex annotation guidelines similar to those of TB-Dense. MATRES refines these guidelines by focusing on a subset of events, annotating only event start times, and reducing the label set to before, after, equal, and vague. These refinements improve inter-annotator agreement and offer a more accessible setting for the task. However, MATRES is not exhaustively annotated. To bridge this gap, we develop OmniTemp, a dataset that adopts the refined MATRES scheme while ensuring complete coverage of all event pairs across entire texts. Further details are provided in §3.
Train Test All
@@
2.2 Zero-Shot Methods
Recent advancements in LLMs offer an opportunity to leverage their vast knowledge for zero-shot approaches Kojima et al. (2022), enabling solutions without training data Zhao et al. (2023). However, few studies have explored LLMs for temporal relation extraction in zero-shot settings. The most notable one is by Yuan et al. (2023), who applied a simple zero-shot chain-of-thought (ZS-CoT) method, where the model is asked about each relation for a given event pair until it answers “yes”. Another effort by Chan et al. (2024) experimented with prompt engineering and in-context learning. Both methods employed a pairwise approach and achieved suboptimal results on the MATRES and TB-Dense datasets. Additionally, the pairwise approach makes these methods cost- and time-inefficient.
3 The OmniTemp Dataset
OmniTemp [^1] Will be released under a custom license that permits free academic use (see Appendix H). is built following the MATRES Ning et al. (2018b) approach; however, instead of annotating events only in consecutive sentences, the annotation is complete, covering all event pairs across the entire document. OmniTemp consists of a set of 30 human-generated English news summaries (Newser.com), derived from the Multi-News dataset Fabbri et al. (2019). We select summaries that portray large events, as these are rich in meaningful event mentions. Each summary contains a set of event mentions, with every pair assigned one of the following relations: before, after, equal, or vague. We now describe OmniTemp’s annotation process (§3.1) along with dataset statistics (§3.2).
3.1 Annotation Process
For the annotation process, we hired three non-expert, native English-speaking annotators (students) to label 30 news summaries (~500 words each) for temporal relations (before, after, equal, or vague) between salient events, following MATRES guidelines and using the EventFull tool Eirew et al. (2025). Starting from ~60 auto-detected event mentions per document, extracted using Cattan et al. (2021), annotators selected 15–18 salient events for full-pair annotation—balancing coverage and annotation quality, as prior work shows agreement declines beyond this range for non-expert annotators Eirew et al. (2025). Final labels were determined by majority vote; in cases of disagreement, the label was set to vague. Further details on the annotation methodology and protocol are provided in Appendix E.
@@ 3.2 Dataset Statistics and Comparison
Table 1 summarizes the OmniTemp dataset’s statistics. Overall, the final annotated version of OmniTemp consists of 30 documents, corresponding to 470 event mentions and 3,483 relations. Appendix J, Table 7 presents the statistics of prominent datasets for the temporal relation extraction task alongside OmniTemp.
@@
Finally, we assess whether transitivity can compensate for the limited annotation scope in datasets like MATRES and TB-Dense, where only consecutive-sentence pairs are annotated. Using the NarrativeTime dataset, we consider only intra- and consecutive-sentence relations, then apply transitive closure to infer additional links. While some long-distance relations are recovered, most inferred relations remain local and sparse. This further highlights the importance of exhaustive annotation. A full analysis is provided in Appendix F.
@@ 4.1 Prompt Structure
Our approach begins with a straightforward yet powerful idea: prompting an LLM to generate the full temporal graph of a document in a single call (Figure 1). The process starts with a general instruction outlining the task. We then employ a two-step procedure, motivated by the observation that directly prompting the model to classify all event pair relations results in a more inconsistent outputs. To address this, and inspired by reasoning-based prompting techniques Wang et al. (2023a); Sun et al. (2024), we first prompt the model ([1] in Figure 1) to construct a free-form timeline that summarizes the temporal flow of the marked events. This primes the model with a broader understanding of event order before making explicit classification decisions. We then instruct the model to predict temporal relations between all event pairs. The input includes the full document with event mentions highlighted using angle brackets and unique identifiers (e.g., <attack(7)>), followed by a list of all possible event pairs. If the list exceeds 200 pairs, it is split into batches and processed separately, then merged in post-processing (see Appendix A).
For the output, we instruct the model to represent relations as a graph, where events serve as nodes and relations as edges, formatted in the DOT language Gansner (2006), which helps suppress free-text explanations and facilitates parsing (an example of the generated timeline is presented in Appendix J, Figure 10).
@@ 4.2 Post Process
LLMs are inherently stochastic and may produce different labels for the same input when run multiple times, leading to unstable outputs, especially for ambiguous event pairs. To address this, inspired by self-consistency methods Wang et al. (2023b) and temporal graph consistency optimization techniques Ning et al. (2018a), we run the model $M=5$ times per document (experimental results show that performance saturates after five generations; see Appendix A) and aggregate the predicted relation labels into a distribution $p_{ij}\in\mathbb{R}^{|\mathcal{R}|}$ for each event pair $(e_{i},e_{j})\in\mathcal{E}\times\mathcal{E}$, where $\mathcal{E}$ denotes the set of events, and $p_{ij}^{r}$ represents the empirical likelihood of label $r\in\mathcal{R}$ across runs. ([2,3] in Figure 1).
We then apply the Integer Linear Programming (ILP) formulation of Ning et al. (2018a) to obtain a globally consistent graph. Specifically, we define binary variables $\mathcal{I}_{r}(i,j)\in\{0,1\}$ for each relation $r$ and pair $(e_{i},e_{j})$, and optimize for enforcing key structural constraints: uniqueness (only one relation per pair), symmetry (e.g., if $r=\texttt{BEFORE}$, then its inverse holds for the reverse pair), and transitivity (e.g., if $A\rightarrow B\rightarrow C$, then $A\rightarrow C$) ([4] in Figure 1). The result is the optimal temporal graph that maximizes model confidence while ensuring global coherence. Further details are provided in Appendix G.
@@ 5 Experimental Setting
Model MATRES TB-Dense NT-6 OmniTemp
@@
Table 2: F1 and Transitive Inconsistency (TI) scores of all models on the four datasets. We use the F1 definition from Ning et al. (2019), and compute the average TI edges per test document, by applying a transitive closure algorithm Warshall (1962) and counting transitive contradictions. (*) For MATRES and TB-Dense using the supervised models, we report results from Tan et al. (2023) as our reproductions were slightly lower; TI is based on our retrained models. Constraints are not reported for these models as they did not improve results. (–) TI is not reported for CoT on MATRES and TB-Dense, as it only predicts gold-labeled pairs and cannot construct a complete graph. Computing TI would require multiple generations over the complete set of pairs, which is prohibitively expensive (see Table 3). Further details are provided in Appendix C.
@@ 5.1 Datasets
In our experiments, we use our own OmniTemp and three additional datasets: MATRES, TB-Dense, and NarrativeTime. [^2] Dataset license details are in Appendix H. Notably, TCR Ning et al. (2018a) and TDD-Manual Naik et al. (2019), two additional datasets for the TRE task, are excluded from our experiments as they omit the vague relation. Since we generate relations for all possible event pairs, the vague label is essential to avoid forcing incorrect relations when context is insufficient. Below, we provide details on the datasets used in our experiments. For our own OmniTemp, we use the first 10 documents as the test set and the remaining documents as the training set, while for all other datasets, we follow their predefined splits.
@@
NT-6. The NarrativeTime (NT) dataset, previously introduced in §2.1, features seven relation types, including the six from TB-Dense and the overlap relation. However, we exclude the overlap relation as it is incompatible for the temporal consistency methods, given that the symmetric counterpart was not annotated. Additionally, NT documents contain an average of 50 events, corresponding to 1,200 relations, per document. Due to LLM context limits, we randomly select 18 events per NT document (see Appendix A).
@@ 5.2 Baseline and State-of-the-Art Models
We compare our zero-shot (GlobalConsistency) method with four models, reproducing state-of-the-art (SOTA) supervised models and a zero-shot chain-of-thought (CoT) baseline method.
@@ 5.3 Ablation Study Design
ZSL-Global. This configuration prompts the LLM once to generate the entire temporal graph directly, ommiting the instruction to generate the timeline of events before classification.
ZSL-Timeline. This configuration includes only the prompting step (with timeline generation) but omits the post-processing step that enforces global consistency.
@@ 6 Results
Table 3: Approximate costs (USD) for the full dataset are shown. For GlobalConsistency, the cost reflects five generations of the complete set of relations. For the CoT method, to reflect a real-world scenario, we generate the complete set of relations once—rather than just the gold ones. Costs are computed using token counts (via OpenAI’s tiktoken) and official model pricing.
@@
Our results are presented in Table 2, with supervised SOTA pairwise models shown in the upper section, and the results of our zero-shot methods—using GPT-4o and DeepSeek-R1—shown in the lower section. [^3] Details on additional models evaluated are provided in Appendix B. Overall, using GPT-4o, our GlobalConsistency approach (§4) outperforms the CoT baseline Yuan et al. (2023) by a large margin across all datasets except TB-Dense. Using DeepSeek-R1, our GlobalConsistency method outperforms the CoT baseline on the densely annotated datasets, NT-6 and OmniTemp, but shows lower performance on the sparsely annotated datasets, MATRES and TB-Dense (further analyzed in §7). However, the improved performance of the CoT method comes at a significant cost, ~7$\times$ more expensive (Table 3), and requires more time, particularly in the DeepSeek-R1 experiments (Table 4). Furthermore, time and cost differences between CoT and our method do not scale linearly with graph size. This is evident in Table 3, where NT-6, with only two more events per document than OmniTemp, incurs higher costs with CoT.
@@
Figure 3: DeepSeek-R1 model performance across different relation subsets: (1) consecutive-sentence event pairs, (2) non-consecutive-sentence event pairs, and (3) full-document event relations. A similar figure for GPT-4o is presented in Figure 6 in Appendix J.
7 Discussion
Event Mentions Count.
We investigate how the number of events in a document impacts the performance of our GlobalConsistency method. Our hypothesis is that models encoding global information are more sensitive to event count, as they must process more information simultaneously. In contrast, pairwise methods, which consider one event pair at a time, are likely less affected. Figure 2 shows MATRES and TB-Dense documents grouped by increasing event counts. [^4] The other datasets we experimented with contain a limited number of events per instance. With the exception of DeepSeek-R1 on MATRES—which demonstrates resilience to large event counts—the results show a performance decline as the number of events increases. This supports our hypothesis and may help explain the performance gap between the CoT method and the ZSL-Global variant in most tests. [^5] In TB-Dense, performance drops sharply for documents with over 25 events. For further analysis, see Appendix C.3.
@@ 7 Discussion
Figure 4: We examine the performance of our prompting method (i.e., ZSL-Timeline) by relation type across two groups of datasets with similar annotation schemes: six-label datasets (TB-Dense and NT-6) and four-label datasets (MATRES and OmniTemp), using DeepSeek-R1. Similar results are observed with GPT-4o, as presented in Figure 7 in Appendix J. The relation labels are: A = after, B = before, I = includes, II = is-included, E = equal, and V = vague.
Event Pair Distance.
We examine whether the annotation distance restriction—where events are annotated only if they are at most one sentence apart, as in MATRES and TB-Dense—can affect model evaluation. To explore this, we evaluate all zero-shot methods on three subsets of OmniTemp and NT-6: the full dataset, event pairs with a sentence distance of at most one (consecutive sentences like in MATRES and TB-Dense), and event pairs with a sentence distance greater than one (non-consecutive sentences). See Figure 3.
@@
To investigate this further, we examine label consistency in documents and event pairs shared between TB-Dense and MATRES, which annotated the same corpus. There are 983 such event pairs. While these datasets follow different annotation guidelines, certain labels should remain consistent. For instance, if an event pair is labeled equal in TB-Dense—indicating that both the start and end times of the two events are the same—then the relation should also be equal in MATRES. Measuring consistency across the four shared relations, we find strong agreement for before and after, with before being the most consistently annotated. However, significant inconsistencies were evident in vague and equal (detailed results are provided in Appendix D). Since in zero-shot settings the model is not trained on a dataset, it does not learn dataset-specific annotation biases. The annotation inconsistency between MATRES and TB-Dense may partly explain the performance drop on these datasets, particularly for vague and equal relations, as well as the lower performance on after compared to before.

### v3

@@ 1 Introduction
The TRE task is formulated as follows: given a pair of event mentions, identify the temporal relation between them (e.g., before, after, equal, include, is included, vague). The task has seen significant progress in recent years with the development of supervised models Tan et al. (2023); Niu et al. (2024). However, these models require large amounts of training data, which is scarce in most domains and languages, and difficult to obtain due to the complexity of manually annotating such relations Pustejovsky and Stubbs (2011).
@@
To address the scarcity of fully annotated datasets, and the inefficiency and global inconsistency of pairwise classification in zero-shot settings, we make the following contributions: [^1] OmniTemp and all experimental code are publicly available at https://github.com/AlonEirew/GlobalZeroShotTRE.
  • We propose a novel zero-shot LLM method that generates the entire temporal graph in a single inference step. Our method prompts the model to produce a free-form summary of the event timeline to guide reasoning or “thinking”, followed by classification of all event pairs, aggregated via a global temporal constraints optimization algorithm to ensure consistency (Figure 1, and §4). This approach significantly reduces computational cost compared to pairwise methods while achieving performance competitive with supervised models (§6).
  • To support global temporal graph extraction, we introduce OmniTemp, a new dataset with exhaustive temporal relation annotations for all event pairs (§3).
  • We discuss how annotation scope and guideline inconsistencies affect zero-shot model assessment. We show that limiting annotations to short-distance event pairs, as well as discrepancies between widely used datasets such as MATRES and TB-Dense, can hinder fair and reliable evaluation of TRE models in zero-shot settings.
@@ 2.1 Temporal Relation Extraction Datasets
The temporal relation extraction task aims to determine the temporal order between pre-extracted events in a text Pustejovsky et al. (2003). For fair and unbiased model evaluation, datasets should provide gold labels for all event pairs or, at a minimum, be randomly sampled from the full set. However, most existing datasets for temporal relation extraction provide only partial annotation due to the complexity and cost of the process Pustejovsky and Stubbs (2011); Naik et al. (2019). As a result, the two most widely used datasets, MATRES Ning et al. (2018b) and TimeBank-Dense (TB-Dense) Chambers et al. (2014), annotate only relations between events in consecutive sentences.
@@
While NarrativeTime provides an exhaustively annotated dataset, it follows complex annotation guidelines similar to those of TB-Dense. MATRES refines these guidelines by focusing on a subset of events, annotating relations only based on event start times, and reducing the label set to before, after, equal, and vague. These refinements improve inter-annotator agreement and offer a more accessible setting for the task. However, MATRES is not exhaustively annotated. To bridge this gap, we develop OmniTemp, a dataset that adopts the refined MATRES scheme while ensuring complete coverage of all event pairs across entire texts. Further details are provided in §3.
2.2 Zero-Shot Methods
Recent advancements in LLMs offer an opportunity to leverage their vast knowledge for zero-shot approaches Kojima et al. (2022), enabling solutions without training data Zhao et al. (2023). However, few studies have explored LLMs for temporal relation extraction in zero-shot settings. The most notable and best-performing approach is by Yuan et al. (2023), who applied a simple zero-shot chain-of-thought (CoT) method. In this method, the model is sequentially asked about each possible relation for a given event pair (e.g., “Is event-a before event-b?”; if “no,” then “Is event-a after event-b?”) until the model answers “yes.” We use Yuan et al. (2023) method as the zero-shot baseline in our experiments. Another effort by Chan et al. (2024) experimented with prompt engineering and in-context learning. Both methods employed a pairwise approach and achieved suboptimal results on the MATRES and TB-Dense datasets. Additionally, the pairwise approach makes these methods cost- and time-inefficient.
3 The OmniTemp Dataset
OmniTemp [^2] Released under a custom license that permits free academic use (see Appendix G). is built following the MATRES Ning et al. (2018b) approach (§2.1); however, instead of annotating events only in consecutive sentences, the annotation is complete, covering all event pairs across the entire document. OmniTemp consists of a set of 30 human-generated English news summaries (Newser.com), derived from the Multi-News dataset Fabbri et al. (2019). We select summaries that depict major events (e.g., a presidential visit abroad, a mass shooting, a major earthquake), as these are typically rich in informative sub-event mentions that describe the event timeline. Each summary contains a set of event mentions, with every pair assigned one of the following relations: before, after, equal, or vague. We now describe OmniTemp’s annotation process (§3.1) along with dataset statistics (§3.2).
3.1 Annotation Process
For the annotation process, we hired three non-expert, native English-speaking annotators (students) to label 30 news summaries (~500 words each) for temporal relations (before, after, equal, or vague) between salient events, following MATRES guidelines and using the EventFull tool Eirew et al. (2025). [^3] The complete annotation guidelines are available within the EventFull annotation tool and GitHub repository https://github.com/AlonEirew/EventFull. Starting from ~60 auto-detected event mentions per document, extracted using Cattan et al. (2021), annotators selected 15–18 salient events for full-pair annotation, balancing coverage and annotation quality, as prior work shows agreement declines beyond this range for non-expert annotators Eirew et al. (2025). Final labels were determined by majority vote; in cases of disagreement, the label was set to vague. Further details on the annotation methodology and protocol are provided in Appendix E.
Train
Test
All
@@ 3.2 Dataset Statistics and Comparison
Table 1 summarizes the OmniTemp dataset’s statistics. Overall, OmniTemp consists of 30 documents, corresponding to 470 event mentions and 3,483 relations. Table 9 in Appendix I, presents the statistics of prominent datasets for the temporal relation extraction task alongside OmniTemp.
@@
Finally, we assess whether transitivity can compensate for the limited annotation scope in datasets like MATRES and TB-Dense, where only consecutive-sentence pairs are annotated. Using the NarrativeTime dataset, we consider only intra- and consecutive-sentence relations, then apply a transitive closure algorithm Warshall (1962) to infer additional links. While some long-distance relations are recovered, most inferred relations remain local and sparse (as illustrated in Figure 11 of Appendix I), further highlighting the importance of exhaustive annotation.
@@ 4.1 Prompt Structure
Our zero-shot approach, referred to as GlobalConsistency, begins with a straightforward yet powerful idea: prompting an LLM to generate the full temporal graph of a document in a single call (Figure 1). The process starts with a general instruction outlining the task. We then employ a two-step procedure, motivated by the observation that directly prompting the model to classify all event pair relations results in a more inconsistent outputs. To address this, and inspired by reasoning-based prompting techniques Wang et al. (2023a); Sun et al. (2024), we first prompt the model ([1] in Figure 1) to construct a free-form timeline that summarizes the temporal flow of the marked events. This primes the model with a broader understanding of event order before making explicit classification decisions. We then instruct the model to predict temporal relations between all event pairs. The input includes the full document with event mentions highlighted using angle brackets and unique identifiers (e.g., <attack(7)>), followed by a list of all possible event pairs.
For the output, we instruct the model to represent relations as a graph, where events serve as nodes and relations as edges, formatted in the DOT language Gansner (2006), which helps suppress free-text explanations and facilitates parsing (an example of the generated timeline is presented in Appendix I, Figure 10).
In documents containing many events that may exceed the model’s input capacity, we generate the complete set of pairs and split them evenly for separate processing. Each split receives the same instructions and the full report with all event mentions marked in it, this is followed by only the relevant subset of event pairs for that split. The predictions are then merged back in post-processing (Further details are provided in Appendix A)
@@ 4.2 Post Process
LLMs are inherently stochastic and may produce different labels for the same input when run multiple times, leading to unstable outputs, especially for ambiguous event pairs. To address this, inspired by self-consistency methods Wang et al. (2023b) and temporal graph consistency optimization techniques Ning et al. (2018a), we run the model $M=5$ times per document, as experimental results show that performance saturates after five generations (see Figure 5 in Appendix A), and aggregate the predicted relation labels into a distribution $p_{ij}\in\mathbb{R}^{|\mathcal{R}|}$ for each event pair $(e_{i},e_{j})\in\mathcal{E}\times\mathcal{E}$, where $\mathcal{E}$ denotes the set of events, and $p_{ij}^{r}$ represents the empirical likelihood of label $r\in\mathcal{R}$ across runs. ([2,3] in Figure 1).
We then apply the Integer Linear Programming (ILP) formulation of Ning et al. (2018a) to obtain a globally consistent graph. Specifically, we define binary variables $\mathcal{I}_{r}(i,j)\in\{0,1\}$ for each relation $r$ and pair $(e_{i},e_{j})$, and optimize for enforcing key structural constraints: uniqueness (only one relation per pair), symmetry (e.g., if $r=\texttt{BEFORE}$, then its inverse holds for the reverse pair), and transitivity (e.g., if $A\rightarrow B\rightarrow C$, then $A\rightarrow C$) ([4] in Figure 1). The result is the optimal temporal graph that maximizes model confidence while ensuring global coherence. Further details are provided in Appendix F.
@@ 5 Experimental Setting
Non-Exhaustive Exhaustive
MATRES TB-Dense NT-6 OmniTemp
@@
Table 2: F1 and Transitive Inconsistency (TI) scores of all models on four datasets, grouped into Non-Exhaustive Annotation (MATRES, TB-Dense) and Exhaustive Annotation (NT-6, OmniTemp). We use the F1 definition from Ning et al. (2019), and compute the average number of TI edges per test document by applying a transitive closure algorithm Warshall (1962) and counting transitive contradictions. (*) For MATRES and TB-Dense with supervised models, we report results from Tan et al. (2023) as our reproductions were slightly lower; TI is based on our retrained models. Constraints are not reported for these models as they did not improve results. (–) TI is not reported for CoT on MATRES and TB-Dense, as it only predicts gold-labeled pairs and cannot construct a complete graph. Computing TI for CoT would require multiple generations over the full set of pairs, which is prohibitively expensive (see Table 3). Further details are provided in Appendix C.
@@ 5.1 Datasets
In our experiments, we use our own OmniTemp and three additional datasets: MATRES, TB-Dense, and NarrativeTime. [^4] Dataset license details are in Appendix G. Notably, TCR Ning et al. (2018a) and TDD-Manual Naik et al. (2019), two additional datasets for the TRE task, are excluded from our experiments as they omit the vague relation. Since we generate relations for all possible event pairs, the vague label is essential to avoid forcing incorrect relations when context is insufficient. Below, we provide details on the datasets used in our experiments. For our own OmniTemp, we use the first 10 documents as the test set and the remaining documents as the training set, while for all other datasets, we follow their predefined splits.
@@
NT-6. The NarrativeTime (NT) dataset, previously introduced in §2.1, features seven relation types, including the six from TB-Dense and the overlap relation. However, we exclude the overlap relation as it is incompatible for the temporal consistency methods, given that the symmetric counterpart was not annotated. Additionally, NT documents contain an average of 50 events, corresponding to 1,200 relations, per document. Due to LLM context limits, we randomly select 18 events per NT document.
@@ 5.2 Baseline and State-of-the-Art Models
We compare our GlobalConsistency method with four models, reproducing state-of-the-art (SOTA) supervised models and a zero-shot chain-of-thought (CoT) baseline method.
@@ 5.3 Ablation Study Design
ZSL-Global. This zero-shot learning (ZSL) configuration prompts the LLM once to generate the entire temporal graph directly, ommiting the instruction to generate the timeline of events before classification.
ZSL-Timeline. This ZSL configuration includes only the prompting step (with timeline generation) but omits the post-processing step that enforces global consistency.
@@ 6 Results
7 Discussion
Event Mentions Count.
We investigate how the number of events in a document impacts the performance of our GlobalConsistency method. Our hypothesis is that models encoding global information are more sensitive to event count, as they must process more information simultaneously. In contrast, pairwise methods, which consider one event pair at a time, are likely less affected. Figure 2 shows MATRES and TB-Dense documents grouped by increasing event counts. [^6] The other datasets we experimented with contain a limited number of events per instance. With the exception of DeepSeek-R1 on MATRES, which demonstrates resilience to large event counts, the results show a performance decline as the number of events increases. This supports our hypothesis and may help explain the performance gap between the CoT method and the ZSL-Global variant in most tests. [^7] In TB-Dense, performance drops sharply for documents with over 25 events. For further analysis, see Appendix C.3.
@@ 7 Discussion
Table 3: Approximate costs (USD) for the full dataset are shown. For GlobalConsistency, the cost reflects five generations of the complete set of relations. For the CoT method, to reflect a real-world scenario, we generate the complete set of relations once—rather than just the gold ones. Costs are computed using token counts (via OpenAI’s tiktoken, and DeepSeek official tokenizer) and official model pricing.
@@
Event Pair Distance.
We examine whether the annotation distance restriction, where events are annotated only if they are at most one sentence apart, as in MATRES and TB-Dense, can affect model evaluation. To explore this, we evaluate all zero-shot methods on three subsets of OmniTemp and NT-6: the full dataset, event pairs with a sentence distance of at most one (consecutive sentences like in MATRES and TB-Dense), and event pairs with a sentence distance greater than one (non-consecutive sentences). See Figure 3.
@@
Figure 3: DeepSeek-R1 model performance across different relation subsets: (1) consecutive-sentence event pairs, (2) non-consecutive-sentence event pairs, and (3) full-document event relations. A similar figure for GPT-4o is presented in Figure 6 in Appendix I.
@@
Figure 4: We examine the performance of our prompting method (i.e., ZSL-Timeline) by relation type across two groups of datasets with similar annotation schemes: six-label datasets (TB-Dense and NT-6) and four-label datasets (MATRES and OmniTemp), using DeepSeek-R1. Similar results are observed with GPT-4o, as presented in Figure 7 in Appendix I. The relation labels are: A = after, B = before, I = includes, II = is-included, E = equal, and V = vague.
@@
To investigate this further, we examine label consistency in documents and event pairs shared between TB-Dense and MATRES, which annotated the same corpus. There are 983 such event pairs. While these datasets follow different annotation guidelines, certain labels should remain consistent. For instance, if an event pair is labeled equal in TB-Dense, indicating that both the start and end times of the two events are the same, then the relation should also be equal in MATRES. Measuring consistency across the four shared relations, we find strong agreement for before and after, with before being the most consistently annotated. However, significant inconsistencies were evident in vague and equal (detailed results are provided in Appendix D). Since in zero-shot settings the model is not trained on a dataset, it does not learn dataset-specific annotation biases. The annotation inconsistency between MATRES and TB-Dense may partly explain the performance drop on these datasets, particularly for vague and equal relations, as well as the lower performance on after compared to before.